OpenStudy (anonymous):

solve for 4: p=4h+4u

OpenStudy (precal):

p=4(h+u) p/(h+u)=4 how odd to ask for solve for 4?

OpenStudy (anonymous):

i mean h lol

OpenStudy (precal):

ok p=4h+4u p-4u=4h (p-4u)/4=h